# Manipulating 3D Molecules in a Fixed-Dimensional SE(3)-Equivariant Latent Space

### v1

@@ A.2 SE(3)-Equivariance Discussion
Let $T\in\mathrm{SE}(3)$ denote a rigid transformation. If the condition nodes are centered at zero and the parameterization $\bm{\Phi}(\bm{\theta},\mathbf{c},t)$ is SE(3)-equivariant, then the likelihood is invariant under $T$:
@@ A.3 Encoder Network
where only the virtual node outputs $[\mathbf{z}_{x},\mathbf{z}_{h}]$ are retained as the final latent code. Since the coordinate updates are designed to be SE(3)-equivariant at each layer, the entire encoder $\bm{\phi}_{\theta}$ is equivariant by construction.
@@ Appendix B VAE Details
We adopt the regularization loss to regularize the latent space. Given the deterministic initial output $[\mathbf{z}_{x},\mathbf{z}_{h}]$, we predict a coordinate-wise Gaussian distribution for each latent dimension:
$\bm{\mu}_{x}=\mathbf{z}_{x},\quad[\bm{\sigma}^{2}_{x},\;\bm{\mu}_{h},\bm{\sigma}^{2}_{h}]=\mathtt{Linear}([\mathbf{z}_{x},\mathbf{z}_{h}]).$ (12)
@@ C.1.2 Inference of BFN
Given prior parameters $\bm{\theta}_{0}$, accuracies $\alpha_{1},\dots,\alpha_{n}$ and corresponding times $t_{i}=i/n$, the $n$-step sampling procedure recursively generates $\bm{\theta}_{1},\dots,\bm{\theta}_{n}$ by sampling $\mathbf{x}^{\prime}$ from $p_{O}(\cdot\mid\bm{\theta}_{i-1},\mathbf{z},t_{i-1})$, $\mathbf{y}$ from $\mathbf{y}\sim p_{R}(\cdot\mid\bm{\theta}_{i-1},\mathbf{z},t_{i-1},\alpha_{i})$, then setting $\bm{\theta}_{i}=h(\bm{\theta}_{i-1},\mathbf{z},\mathbf{y})$. And pass the Neural Network. The final sample is drawn from $p_{O}(\cdot\mid\bm{\theta}_{n},\mathbf{z},1)$.
@@ C.2 MolFLAE Reconstruction Loss
SE(3)-equivariant encoding:
  (_$\bigl{(}\_,[\mathbf{z}_{x},\mathbf{z}_{h}]\bigr{)}\leftarrow\bm{\phi}_{\theta}\!\bigl{(}(\mathcal{M},\mathcal{Z})\bigr{)}$
@@
  $[\bm{\mu}_{x},\bm{\sigma}^{2}_{x},\bm{\mu}_{h},\bm{\sigma}^{2}_{h}]\leftarrow\texttt{Linear}\bigl{(}[\mathbf{z}_{x},\mathbf{z}_{h}]\bigr{)}$
@@ C.4 MolFLAE Decoding Process
Input: Network $\bm{\Phi}$, latent code $\mathbf{z}\in\mathbb{R}^{N_{Z}(3+D_{Z})}$, total steps $N$, number of atoms $N_{M}$, number of types $D_{M}$, noise levels $\sigma_{1}$, $\beta_{1}$
Output: Sampled molecule $[\hat{\mathbf{x}},\hat{\mathbf{v}}]$
@@
return $[\hat{\mathbf{x}},\hat{\mathbf{v}}]$
@@ Appendix D Molecules Property Metrics
In our interpolation experiments, we employed the following metrics to fully assess the chemical and physical properties of the molecules:
@@
Similarity $\text{Similarity Preference}=\frac{S_{t}-S_{s}}{S_{t}+S_{s}}$
@@
Appendix E Ablation Study
@@
Appendix F Hyper-parameter Settings
@@
Appendix G Broader Impacts and Safety Discussion

### v2

@@ A.2 E(3)-Equivariance Discussion
Let $T\in\mathrm{E}(3)$ denote a rigid transformation. If the condition nodes are centered at zero and the parameterization $\bm{\Phi}(\bm{\theta},\mathbf{c},t)$ is E(3)-equivariant, then the likelihood is invariant under $T$:
@@ A.3 Encoder Network
where only the virtual node outputs $\mathbf{z}_{x},\mathbf{z}_{h}]$ are retained as the final latent code. Since the coordinate updates are designed to be E(3)-equivariant at each layer, the entire encoder $\bm{\phi}_{\theta}$ is equivariant by construction.
@@ Appendix B VAE Details
We adopt the regularization loss to regularize the latent space. Given the deterministic initial output $\mathbf{z}_{x},\mathbf{z}_{h}]$, we predict a coordinate-wise Gaussian distribution for each latent dimension:
$\bm{\mu}_{x}=\mathbf{z}_{x},\quad[\bm{\sigma}^{2}_{x},\;\bm{\mu}_{h},\bm{\sigma}^{2}_{h}]=\mathtt{Linear}(\mathbf{z}_{h}).$ (12)
@@ C.1.2 Inference of BFN
Given prior parameters $\bm{\theta}_{0}$, accuracies $\alpha_{1},\dots,\alpha_{n}$ and corresponding times $t_{i}=i/n$, the $n$-step sampling procedure recursively generates $\bm{\theta}_{1},\dots,\bm{\theta}_{n}$ by sampling $\mathbf{x}^{\prime}$ from $p_{O}(\cdot\mid\bm{\theta}_{i-1},\mathbf{z},t_{i-1})$, $\mathbf{y}$ from $\mathbf{y}\sim p_{R}(\cdot\mid\bm{\theta}_{i-1},\mathbf{z},t_{i-1},\alpha_{i})$, then setting $\bm{\theta}_{i}=h(\bm{\theta}_{i-1},\mathbf{z},\mathbf{y})$, and pass the result to the neural network. The final sample is drawn from $p_{O}(\cdot\mid\bm{\theta}_{n},\mathbf{z},1)$.
@@ C.2 MolFLAE Reconstruction Loss
E(3)-equivariant encoding:
  $\bigl(\_,[\mathbf{z}_{x},\mathbf{z}_{h}]\bigr)\leftarrow\bm{\phi}_{\theta}\!\bigl((\mathcal{M},\mathcal{Z})\bigr)$
@@
  $\bm{\mu}_{x}\leftarrow\mathbf{z}_{x},\quad[\bm{\sigma}^{2}_{x},\;\bm{\mu}_{h},\bm{\sigma}^{2}_{h}]\leftarrow\mathtt{Linear}(\mathbf{z}_{h})$
@@ C.4 MolFLAE Decoding Process
Input: Network $\bm{\Phi}$, latent code $\mathbf{z}\in\mathbb{R}^{N_{Z}(3+D_{Z})}$, total steps NN, number of atoms $N_{M}$, number of types $D_{M}$, noise levels $\sigma_{1}$, $\beta_{1}$
Output: Sampled molecule $\hat{\mathbf{x}},\hat{\mathbf{v}}$]
@@
return $\hat{\mathbf{x}},\hat{\mathbf{v}}$]
@@ Appendix D Molecules Property Metrics
Following the setup of EDM[15], in our unconditional generation experiments, we employed the following metrics to evaluate the quality of generated molecules:
Atom Stability: Proportion of atoms with valid bond counts.
Molecular Stability: Proportion of molecules where all atoms are stable.
Validity: Proportion of molecules with RDKit-parsable SMILES.
Novelty: Proportion of molecules whose SMILES are not in the training set.
Uniqueness: Proportion of unique molecules in the generated set.
In our interpolation experiments, to fully assess the chemical and physical properties of intermediary molecules generated during interpolation, we employed the following metrics:
@@
$\text{Similarity Preference}=\frac{S_{t}-S_{s}}{S_{t}+S_{s}}$
@@
Appendix E Results on drug-likeness metrics on GEOM-Drugs
To provide a more comprehensive evaluation, we conduct experiments on the GEOM-Drugs dataset with additional drug likeness metrics, where MolFLAE showed significant improvements over existing baselines. The results are summarized in Table 7. It should be noted that we were unable to locate open resources for GeoBFN’s checkpoints on GEOM-Drugs, so experiments on GeoBFN were not conducted.
In details, we sample 10,000 molecules from each model with their default settings, obtaining atom positions and types, and then inferred bond types using OpenBabel. We then fix the bond order using Schordinger due to some bugs in OpenBabel. The final molecules are then be evaluated using RDKit for the following metrics: QED, SA, Lipinski, and Strain Energy.
Table 7: Comparison of models on QED, SA, Lipinski and Strain Energy on GEOM-Drugs dataset.
# Metrics QED (↑) SA (↑) Lipinski (↑) Strain Energy (↓)
Data 0.64 0.84 4.80 80.19
EDM 0.36 0.59 4.26 705.2
GeoLDM 0.40 0.63 4.31 446.1
UniGEM 0.36 0.63 4.24 490.4
MolFLAE 0.60 0.75 4.75 84.66
Appendix F Ablation Study
@@
Moreover, as our iterative decoding process is a little complex, we tried to simplify it to a one-step decoding variant. However, this approach failed to generate any valid molecules.
Appendix G Hyper-parameter Settings
@@
Appendix H Broader Impacts and Safety Discussion